OpenStudy (anonymous):

Find the domain and range of F(x) = -2 (x-1)^2 +7

OpenStudy (anonymous):

\((x-1)^2\geq 0\) because it is a square so cannot be negative therefore \(-2(x-1)^2\leq 0\) because of the minus sign and therefore \[-2(x-1)^2+7\leq 7\]

OpenStudy (anonymous):

that tells you the range is \((-\infty, 7]\) domain is all real numbers because it is a polynomial

OpenStudy (anonymous):

Domain: all real numbers Range: \[y \le 7\]

OpenStudy (anonymous):

satellite got it right

OpenStudy (anonymous):

Thanks C: